OpenStudy (anonymous):

Use Chain Rule to Solve F(x)=f(f(x) and G(x)=(F(x))^2. You also know that f(3)=8 f(8)=3 f'(8)=11 f'(3)=4. Find F'(3) and G'(3) I have been stuck on this for hours.

OpenStudy (anonymous):

differentiate F(x)= f(f(x))

OpenStudy (anonymous):

F'(x)= f'(f(x)) f'(x)

OpenStudy (anonymous):

now at x=3 F'(3)= f'(f(3)) f'(3)

OpenStudy (anonymous):

So would it be 4*3*3?

OpenStudy (anonymous):

F'(3)= f'(f(3)) f'(3) = f'(8)4= 11*4=44 do the same for other one

OpenStudy (anonymous):

fine?

OpenStudy (anonymous):

Still a little lost on the G(x) one

OpenStudy (anonymous):

So would it be 2(44)*11?

OpenStudy (anonymous):

G(x)=(F(x))^2 G'(x)= 2 F(x)* F'(x) at x=3 G'(3)= 2 F(3)* F'(3) u know F'(3) =44 from previous question so u have to find F(3) only

OpenStudy (anonymous):

use this for getting F(3) F(x)=f(f(x)

OpenStudy (anonymous):

does it help u?

OpenStudy (anonymous):

So wait is it 44*44?

OpenStudy (anonymous):

F(3)=f(f(3)) = f(8) =3 so G'(3)= 2 F(3)* F'(3)= 2 *3*44

OpenStudy (anonymous):

I get it now thank you so much!

OpenStudy (anonymous):

:)